# MyFramelinecolor=black, outerlinewidth=2pt, roundcorner=20pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=codeBackground

### v2

@@ 1 Introduction
In the same post, Buterin adds that the challenge of obtaining a proof of useful work is “probably not feasible”. The transition of Etherium from PoW to PoS is, to a large extent, attributed to this impossibility conjecture.
@@ 2 Overview of our PoUW
      Compute $z=\mathcal{\mathcal{missing}}O(C^{\prime})$.
@@
      Compute $z=\mathcal{\mathcal{missing}}O(\mathsf{tr})$.
@@ 5 Definitional Framework for Proofs of Useful Work
Hardness: For any $\delta>0$ and any algorithm $\mathsf{Solve}^{\star}$ with unbounded polynomial-time preprocessing and whose online running time is $(\alpha(t(n)))^{1-\delta}$, it holds that
$\Pr\left[\mathsf{Verify}(\sigma,\pi)=1\right]\leq\epsilon(n),$
where the probability is over the randomness of $\sigma\leftarrow\{0,1\}^{\lambda}$ sampled independently uniformly at random, computing $\pi\leftarrow\mathsf{Solve}^{\star}(\sigma)$, and over the random oracle.
@@ A PoUW for $f$, denoted $(\mathsf{Solve},\mathsf{Verify})$ is said to be optimal if
  • $\alpha(t(n))=t(n)\cdot(1+o(1))$ and $\beta(t(n))=t(n)\cdot(1-\omega(1))$
@@
Remark 5.6 (On hardness).
Hardness considers a scenario where an attacker gets a fresh challenge of the PoUW. We provide the prover with a complexity budget sufficient to run the honest prover and ask it to perform significantly better, i.e., have much better than $\epsilon$ probability of causing the verifier to accept. In words, this should be interpreted as saying that the honest provers strategy is optimal, up to small terms.
One can introduce stronger notions of security, considering the success probability of a malicious prover when working on multiple instances in parallel (i.e., whether amortization exists), or when the adversary receives challenges in a streaming fashion and it only needs to show slight advantage over the naive process. The latter property is essentially what is needed for Bitcoin-style blockchains, and we formalize the security definition in Appendix B We conjecture that our PoW mechanism will satisfy all properties satisfied by the PoW of Bitcoin, and focus in the main body on the simpler definition to simplify presentation and convey the main new ideas.
@@ 5.1 A (Trivial) PoUW for Any Function
  • $\mathsf{Solve}$: on input $\sigma$ and $\mathsf{x}$, the prover solves the task by running Alg on $\mathsf{x}$ to get $\mathsf{y}$. Also, independently, it does a proof of work to convince the verifier that it spent enough time. Specifically, it chooses $T=c\cdot t(n)$ nonces $\omega_{1},\ldots,\omega_{T}$ for the random oracle, and computes $z_{i}=\mathcal{O}(\sigma,\omega_{i})$. It then sets $\pi$ to be the $\omega_{i}$ for which $z_{i}$ starts with the most 00’s.
@@ 6 A PoUW for Matrix Multiplication
This section is structured as follows. We first present the canonical matrix multiplication algorithm that we assume (referred to as MatMul). Our protocols are roughly based on the random self reducibility property of matrix multiplication. Our first protocol is rather straightforward, and while being better than the generic one in Section 5.1, it is still not an optimal PoUW. Our second protocol combines the idea of random self reducibility of matrix multiplication together with an additional structural property of the algorithm we us. This protocol, as we show, is an optimal PoUW under a new algorithmic conjecture.
@@ 6.5 A Second Instantiation
Recall that the algorithm that we described in Section 6.1 can be used to multiply any two matrices in $O(n^{\omega_{r}+1}/r)$ time, and in particular, to multiply random rank $r$ ones. One can wonder if there is a way to get a speedup by utilizing the low rank property. Suppose that the matrices one wants to multiply are denoted $A$ and $B$, and that the low rank decomposition of them is $A=A_{L}\cdot A_{R}$ and $B=B_{L}\cdot B_{R}$, where $A_{L},B_{L}$ are $n\times r$ and $A_{L},B_{L}$ are $r\times n$. Now, one can generate all the intermediates by computing a partial sum $T_{z}=\sum_{i=1}^{z}A_{R}\cdot B_{L}$ in blocks of size $r$ (so $z$ ranges in $1,\ldots,n/r$), and then multiplying $A_{L}\cdot T_{z}\cdot B_{R}$. Each of the latter can be done by fast rectangular matrix multiplication, costing $n^{\omega_{r}}$, but there are $n/r$ different values of $z$. So, overall, we obtain the same running time of $O(n^{\omega_{r}+1}/r)$, as the general purpose algorithm.
@@ 7 Using Native AI Workloads for Blockchain Consensus
Utilizing the structure and predictability of MatMuls in AI workloads can provide significant dpeedups in practice (i.e., reduction in overhead). We briefly discuss two ideas. First, in several applications, notably inference, the weight matrices are known and are public at the start of the process, and so their hash can be preprocessed, avoiding the need to hash them at every layer “on the fly.” Second, in training applications, the matrices in a given level are typically the output of a product from the previous level (after activations). Thus, in principle, one need not commit to the matrices from scratch in every level, but rather one can utilize a commitment to the very first level and then “recompute” the needed matrices.
@@
The computational burden in AI models primarily arises from matrix multiplications. For matrices of size $n\times n$ and $n\times m$, the multiplication has a complexity of $nnm$. When $n=n=m$, this simplifies to $O(n^{3})$. Other operations, including activation functions, normalizations, rounding, etc. typically scale with $O(n^{2})$ as they operate element-wise.

### v3

@@ 1 Introduction
In the same post, Buterin adds that the challenge of obtaining a proof of useful work is “probably not feasible”. The transition of Ethereum from PoW to PoS is, to a large extent, attributed to this impossibility conjecture.
@@ 2 Overview of our PoUW
      Compute $z=\mathcal{O}(C^{\prime})$.
@@
      Compute $z=\mathcal{O}(\mathsf{tr})$.
@@ 5 Definitional Framework for Proofs of Useful Work
Hardness: For any $\delta>0$ and any algorithm $\mathsf{Solve}^{\star}$ with polynomial-time preprocessing and whose online running time is $(\alpha(t(n)))^{1-\delta}$, such that
$\Pr\left[\mathsf{Verify}(\sigma,\pi)=1\right]<\epsilon(n),$
where the probability is over the randomness of $\sigma\leftarrow\{0,1\}^{\lambda}$ sampled uniformly at random, computing $\pi\leftarrow\mathsf{Solve}^{\star}(\sigma)$, and over the random oracle.
In the above we give the adversary a smaller compute budget and assume that it cannot perform as well as the honest solver. There are many alternative ways to define hardness, depending on how flexible or strict one wants to be. For example, one could be more nuanced about the ratio between the work and the probability of verification succeeding. This is, in fact, more aligned with the typical way proofs of work are used in practice: if one invests twice more work, then its reward is factor two higher. The hardness property in this case can be formalized by parametrized the running time of the attacker by $\gamma(\alpha(n))$ and saying that $\Pr\left[\mathsf{Verify}(\sigma,\pi)=1\right]\leq(\gamma(\alpha(n))/\alpha(n))\cdot\epsilon(n)$. For example, if the adversary invests half of the resources the honest solver does, then its expected reward is exactly half.
Additionally, one can introduce stronger notions of security, considering the success probability of a malicious prover when working on multiple instances in parallel (i.e., whether amortization exists), or when the adversary receives challenges in a streaming fashion and it only needs to show slight advantage over the naive process. The latter property is essentially what is needed for Bitcoin-style blockchains, and we formalize the security definition in Appendix B. We conjecture that our PoW mechanism will satisfy all properties satisfied by the PoW of Bitcoin, and focus in the main body on the simpler definition to simplify presentation and convey the main new ideas.
@@ A PoUW for $f$, denoted $(\mathsf{Solve},\mathsf{Verify})$ is said to be optimal if
  • $\alpha(t(n))=t(n)\cdot(1+o(1))$ and $\beta(t(n))=t(n)\cdot(1-1/c)$ for $c>0$.
@@ 5.1 A (Trivial) PoUW for Any Function
  • $\mathsf{Solve}$: on input $\sigma$ and $\mathsf{x}$, the prover solves the task by running Alg on $\mathsf{x}$ to get $\mathsf{y}$. Also, independently, it does a proof of work to convince the verifier that it spent enough time. Specifically, for a constant $c$ or perhaps slightly super-constant (see discussion below), it chooses $T=c\cdot t(n)$ nonces $\omega_{1},\ldots,\omega_{T}$ for the random oracle, and computes $z_{i}=\mathcal{O}(\sigma,\omega_{i})$. It then sets $\pi$ to be the $\omega_{i}$ for which $z_{i}$ starts with the most 00’s.
@@ 6 A PoUW for Matrix Multiplication
This section is structured as follows. We first present the canonical matrix multiplication algorithm that we assume (referred to as MatMul). Our protocols are roughly based on the random self reducibility property of matrix multiplication. Our first protocol is rather straightforward, and while being better than the generic one in Section 5.1, it is still not an optimal PoUW. Our second protocol combines the idea of random self reducibility of matrix multiplication together with an additional structural property of the algorithm we use. This protocol, as we show, is an optimal PoUW under a new algorithmic conjecture.
@@ 6.5 A Second Instantiation
Recall that the algorithm that we described in Section 6.1 can be used to multiply any two matrices in $O(n^{\omega_{r}+1}/r)$ time, and in particular, to multiply random rank $r$ ones. One can wonder if there is a way to get a speedup by utilizing the low rank property. Suppose that the matrices one wants to multiply are denoted $A$ and $B$, and that the low rank decomposition of them is $A=A_{L}\cdot A_{R}$ and $B=B_{L}\cdot B_{R}$, where $A_{L},B_{L}$ are $n\times r$ and $A_{L},B_{L}$ are $r\times n$. Now, one can generate all the intermediates by computing a partial sum $T_{z}=\sum_{i=1}^{z}A_{R}\cdot B_{L}$ in blocks of size $r$ (so $z$ ranges in $1,\ldots,n/r$), and then multiplying $A_{L}\cdot T_{z}\cdot B_{R}$. Each of the latter can be done by fast rectangular matrix multiplication, costing $n^{\omega_{r}}$, but there are $n/r$ different values of $z$. So, overall, we obtain the same running time of $O(n^{\omega_{r}+1}/r)$, as the general purpose algorithm. Using the state of the art rectangular fast matrix multiplication, we can see that the above assumption is true (because the number bits needed to write down all intermediates – $\Theta(n^{3}/r)$ – is always a lower bound on the running time and $\omega_{r}=2$ for sufficiently small $r$.
@@ 7 Using Native AI Workloads for Blockchain Consensus
Utilizing the structure and predictability of MatMuls in AI workloads can provide significant speedups in practice (i.e., reduction in overhead). We briefly discuss two ideas. First, in several applications, notably inference, the weight matrices are known and are public at the start of the process, and so their hash can be preprocessed, avoiding the need to hash them at every layer “on the fly.” Second, in training applications, the matrices in a given level are typically the output of a product from the previous level (after activations). Thus, in principle, one need not commit to the matrices from scratch in every level, but rather one can utilize a commitment to the very first level and then “recompute” the needed matrices.
@@
The computational burden in AI models primarily arises from matrix multiplications. For matrices of size $n\times k$ and $k\times m$, the multiplication has a complexity of $nkm$. When $n=k=m$, this simplifies to $O(n^{3})$. Other operations, including activation functions, normalizations, rounding, etc. typically scale with $O(n^{2})$ as they operate element-wise.
@@
Acknowledgements
We thank Yonatan Sompolinsky for multiple discussions on the topic of this work and for valuable comments on the manuscript. We thank Ohad Klein for valuable feedback and particularly for suggesting the FMM-based algorithm for computing the intermediates. We thank Eylon Yogev and Mark Zhandry for useful remarks and suggestions on this manuscript.